# Artificial Intelligence HCI Human-Computer Interaction

### v1

@@ 4.2. RQ2: Participants’ Concerns and Conditions for Integrating AI Support
Almost all participants were generally open to using or experimenting with an AI tool in their QDA process. However, most of these participants emphasized that certain conditions should be met or that they would still want to have full control over the QDA process. While a few participants were eager to integrate AI into their workflow, one was strongly opposed to the use of AI, stating that they did not see any useful applications for it in their workflow.
Several participants recognized the potential of AI to automate repetitive tasks – such as data preprocessing, organizing raw data, or annotating thousands of survey responses. Other participants saw an opportunity for AI to contribute ideas and enrich the analytical process with new views and perspectives. They expressed interest in AI’s ability to identify relationships and suggest novel patterns in the data that might be overlooked, complementing human analysis. However, they also expressed various concerns, which we categorized into four main themes: research ethics, personal involvement and responsibility, quality of AI output, and specific requirements of AI tools. Interestingly, our findings revealed no correlation between participants’ level of experience in QDA and their willingness to adopt AI or their concerns regarding its integration. Both novice and experienced researchers expressed openness to exploring AI tools while emphasizing the importance of safeguards and customization to align with their workflows.
@@ 4.2.2. Personal Involvement and Responsibility
Participants were reluctant to become overly dependent on AI tools, emphasizing the need to remain directly connected with their data and engaged with the analysis. Even with AI assistance, researchers must retain responsibility for their research outcomes. This sentiment reflects a concern for preserving the integrity and independence of academic research, ensuring that AI serves as a supportive tool rather than replacing the researcher’s role. Moreover, conducting QDA involves learning from the process and improving from experiences. As one researcher explained, “doing the coding has helped me immensely with the analysis, I think it is necessary and I would not want to give that up.” Participants agreed that AI tools should enhance human capabilities rather than degrade them or make the process more cumbersome. On a broader level, one participant voiced concerns about the potential for AI to replace human coders, resulting in the displacement of jobs within the research community.
@@ 4.2.4. Conditions and Requirements
Given their concerns, participants formulated conditions and requirements for the potential use of AI in QDA, emphasizing tools needed to be transparent, customizable, and supportive of human oversight. To meet privacy considerations, a few participants stressed that AI models should be able to operate in offline mode, ensuring that no data would be uploaded or shared with third parties. Addressing the need for personal involvement and responsibility, several participants emphasized the importance of evaluating the performance and reliability of AI tools before fully integrating them into their workflows. Many participants were particularly concerned about retaining control and agency over the analysis process – with several participants stressing the ability to adjust the level of AI’s influence, preferring a more interactive collaboration with AI rather than a fully automated analysis. A majority of participants also expressed a strong desire for mechanisms to check and verify AI outputs, ensuring the results could be trusted. Several participants further wanted explanations for AI-generated codes: “I would want an explanation for each code, or at least an example from the data to show where the decision was coming from.” Other participants wanted to take random samples of AI outputs and check them or compare them to their own results. Generally, many participants expressed a preference for an interactive approach to AI integration, where they could collaborate with the AI rather than relying on it entirely. Some participants also wanted the ability to tinker with the AI’s settings, modify the model’s parameters, and tailor the outputs to better suit their research needs.
@@
Participants were generally open to using AI tools in their QDA processes, provided that these tools enhance rather than replace their skills. By delegating time-consuming tasks to AI, researchers could focus on deeper analysis and uncover insights that require human interpretation and critical thinking. Others saw an opportunity in AI to enrich the analysis with new views and perspectives. Hence, AI was perceived as both: a tool for efficiency and a catalyst for analysis, with the potential to benefit the depth and breadth of QDA. Participants expressed several concerns and mentioned conditions that had to be met. They suggested developing AI tools that operate offline to allow for greater control over the data’s use and storage. Quality and trustworthiness of AI outputs were another major concern, especially given AI’s current limitations in dealing with nuanced data and the risk of generating false information. Participants stressed the importance of rigorously testing and validating AI tools before integrating them into their research workflows to ensure their reliability in research contexts. Still, they saw potential in AI tools that are developed in alignment with their needs, with a focus on transparency, control, and ethical compliance. Notably, our findings revealed no clear correlation between participants’ level of experience in QDA and their willingness to adopt AI or their concerns. This underscores that the identified needs and conditions for AI integration are shared across different levels of expertise.
@@ 5. A Framework for AI-based Support that Aligns with HCI Researchers’ Real-Life QDA Workflows
Figure 2. Our framework for integrating AI into QDA, depicting 3 stages from less to more AI involvement, visualizing possible roles for AI we deduced from the data for each stage.
[…]
Figure 2 shows our framework, visualizing less to more AI involvement in 3 stages with a widening cone. The different stages contain visualizations for the relationship between AI and human researchers. Stage 1: AI does not directly impact human workflow. The figure representing a person is larger than the figure representing a robot. Stage 2: AI collaborates with human. The figure representing a person is the same size as the figure representing a robot. Stage 3: AI is the main driver. The figure representing a person is smaller than the figure representing a robot. Pictograms visualize 16 different possible roles for AI in the 3 stages. A pictogram of a robot represents AI, and different pictograms visualize the different tasks. For example, a pictogram of a robot and 2 fighting persons is shown next to AI as a mediator.
@@ 5.1. Minimal AI Involvement: Supporting Human-Centric Workflows
[…]
@@
[…]
@@
[…]
@@ 5.2. Moderate AI Involvement: Human-AI Collaboration
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 5.3. High AI Involvement: AI-Led Workflows
[…]
@@
[…]
@@
[…]
@@ 6. Discussion
6.1. Situating Our Findings in Prior Work
Our work complements the numerous studies that have explored the use of AI assistance in QDA (Feuston and Brubaker, 2021; Jiang et al., 2021; Marathe and Toyama, 2018; Gao et al., 2023; Kapania et al., 2024; Zhang et al., 2024) by offering a more granular view of how AI might align with researchers’ needs and practices, in the form of detailed interviews with HCI researchers to understand their workflows. Previous studies have highlighted common challenges in QDA, such as the messiness and uncertainty of data analysis (Jiang et al., 2021), which our participants also acknowledged. However, our interviews reveal that not all pain points are areas where researchers desire automation; for example, participants often prefer to engage directly with the “messiness” of their data to better understand it. Consistent with Jiang et al. (Jiang et al., 2021), our participants emphasized the intrinsic value they find in QDA despite its resource-intensive nature, such as the satisfaction of uncovering patterns and the learnings that come from collaboration. This reflects the dual nature of AI adoption in research, where there is both enthusiasm for its potential benefits and caution regarding what might be lost in the process (Feuston and Brubaker, 2021). However, our study identifies a notable shift in attitude with the proliferation of AI tools: some researchers are more open to AI collaboration, especially in roles that allow them to retain agency, such as providing suggestions or validation.
The diversity of workflows, concerns, and levels of willingness highlighted in our study underscores the need for a nuanced framework for AI integration in QDA. Previous studies have explored various applications for AI within QDA and proposed broad categories for AI involvement, providing valuable foundational insights (Feuston and Brubaker, 2021). Building on this work, our study offers a more detailed perspective by proposing different roles for AI at increasing levels of involvement, tailored to the specific needs of researchers performing QDA. Jiang et al. (Jiang et al., 2021) applied the AI task delegability framework by Lubars et al. (Lubars and Tan, 2019) to assess which QDA tasks AI should perform, considering factors like motivation, risk, trust, and difficulty. Similarly, Eschrich et al. (Eschrich and Sterman, 2024) propose a framework for using LLMs in QDA, focusing on how they propose or refute codes and how humans verify AI outputs. Our framework extends these approaches by ideating tangible scenarios for AI support that go beyond mere coding, such as preprocessing data, onboarding new researchers, mediating conflicts, and acting as a conversational partner. It accommodates varying levels of AI involvement while retaining human agency, providing flexible guidance that aligns with real-world researcher needs and preferences. We suggest that our framework should be viewed as inspirational rather than prescriptive, recognizing that many factors must be considered when integrating AI into qualitative research practices.
@@ 6.2. Putting Our Framework Into Practice
Our framework aims to empower those who want to incorporate AI tools into QDA, but are unsure of how to approach it or where to intervene. A naive approach is to jump towards AI doing the qualitative coding itself, but this was a contentious topic for our participants. Instead, our framework offers numerous ways AI tools can be incorporated in ways that facilitate QDA without compromising research autonomy.
Researchers who are curious about incorporating AI tools such as LLMs but are worried about how this may taint the QDA process can strive for minimal AI involvement, such as AI as a training tool. For companies developing QDA software (e. g., nVivo or MAXQDA), many of which are already adding AI tools to their software, our framework hopes to inspire additional insights into the numerous ways AI tools can be developed to support QDA. Our framework presents three levels of AI integration in QDA workflows, from minimal involvement to full autonomy. Each level identifies specific roles AI can play, ranging from technical support to autonomous decision-making. Determining the appropriate level and role requires careful consideration of factors like research goals, data types, and QDA methods. Below, we outline key considerations for integrating AI into QDA.
@@
Related work shows that coding performance can depend on various factors of complexity, e. g., the required level of judgment and interpretation  (Ortloff et al., 2023; Kirsten et al., 2024). While higher task complexity is an indicator of reduced AI involvement, routine and repetitive tasks like transcription offer the potential for AI leadership, freeing researchers to focus on more complex interpretive tasks. AI should be employed where its capabilities align with the complexity of the task. For example, when the model shows high certainty in a task, it might take a more leading role, while in tasks involving ambiguity or high interpretive judgment, AI could step back. In addition, researchers’ comfort and familiarity with AI tools should guide their level of involvement (Lubars and Tan, 2019). Recent work on algorithmic resignation (Bhatt and Sargeant, 2024) discusses how restricting AI outputs or displaying performance disclaimers might aid the effective use of AI.
@@
Integrating AI into QDA also involves evaluating ethical considerations, such as data privacy and security, to ensure compliance with ethical and legal standards and regulations. Additionally, participants in human subject studies should be informed during the consent process if AI will be used to analyze their data. Further, the harms of engaging with AI outputs need to be discussed, and safeguards should be applied when feeding data into AI systems. Finally, depending on the level of AI autonomy, a normative discussion on violations of intellectual integrity is necessary. For example, it may be necessary to report the level of AI usage in research outputs to maintain transparency and accountability in the research process. Researchers already use AI throughout their research process (e. g. (Kapania et al., 2024; Argyle et al., 2023; Hämäläinen et al., 2023; Oppenlaender and Hämäläinen, 2023)), and popular QDA software increasingly integrates AI features – calling the scientific community, and specifically the HCI community working at the intersection of humans and technology, to engage this debate and establish guidelines and best practices, e.g., on transparency. Kapania et al. (Kapania et al., 2024) discuss the ethical guidelines for navigating these risks and concerns.
These decision points highlight the need for AI tools that are flexible, transparent, and capable of providing different levels of involvement based on the specific research context. Future tool development should focus on creating AI systems that allow researchers to make informed decisions at each point, providing clear options for customizing AI use, ensuring data privacy, and maintaining ethical standards.
@@ Appendix B Complete Codebooks from Our Interviews
$\hookrightarrow$ Explainability AI have an explanation for its response “Can they provide justification/explanation?” 7
@@ Appendix C Examples of Participants’ Coding Workflow Drawings
Figure 15. Four examples of coding workflow drawings created by our participants during interviews. The drawings varied considerably in the level of detail they provided. Some included in-depth descriptions of specific actions during each step, while others focused more on visuals, providing details of the process only verbally.
[…]
This figure shows four different photos of workflow drawings created by participants during the interviews. One workflow is very complex, containing many substeps. The second workflow includes only three main steps (Interviews, followed by coding using tool support, and finally, the paper itself). The participant wrote short notes, such as ”intercoder agreement” and ”thematic analysis,” next to the arrows. The third workflow contains only icons and no text. It depicts a six-step process: 1) a piece of paper and a broom, 2) a piece of paper, a pen, and eyes, 3) two speech bubbles, 4) a mind map, 5) three people, and 6) a piece of paper labeled ”Results.” The last picture is simpler; it contains only a circle consisting of four steps: Collect data, meet and discuss, code, and iterate codebook. Below this, the participant created a small table indicating the columns their typical codebook would have.

### v2

@@ 4.2. RQ2: Participants’ Concerns and Conditions for Integrating AI Support
Almost all participants were open to exploring AI tools in their QDA process. However, most of these participants emphasized that certain conditions should be met or that they would still want to have full control over the QDA process. While a few participants were eager to integrate AI into their workflow, one was strongly opposed to the use of AI, stating that they did not see any useful applications for it in their workflow.
Several participants recognized the potential of AI to automate repetitive tasks – such as data preprocessing, organizing raw data, or annotating thousands of survey responses. Others saw an opportunity for AI to contribute ideas and enrich the analytical process with new views and perspectives. They expressed interest in AI’s ability to identify relationships and suggest novel patterns in the data that might be overlooked, complementing human analysis. However, they also expressed various concerns, which we categorized into four main themes: research ethics, personal involvement and responsibility, quality of AI output, and specific requirements of AI tools. We found no correlation between participants’ level of experience in QDA and their willingness to adopt AI or their concerns regarding its integration. Both novice and experienced researchers expressed openness to exploring AI tools while emphasizing the importance of safeguards and customization to align with their workflows.
@@ 4.2.2. Personal Involvement and Responsibility
Participants were reluctant to become overly dependent on AI tools, emphasizing the need to remain directly connected with their data and engaged with the analysis. Even with AI assistance, researchers must retain responsibility for their research outcomes. This sentiment reflects a concern for preserving the integrity and independence of academic research, ensuring that AI serves as a supportive tool rather than replacing the researcher’s role. Moreover, conducting QDA involves learning from the process and improving from experiences. As one researcher explained, “doing the coding has helped me immensely with the analysis, I think it is necessary and I would not want to give that up”. Participants agreed that AI tools should enhance human capabilities rather than degrade them or make the process more cumbersome. One participant expressed concerns about job displacement, cautioning against AI fully replacing human coders.
@@ 4.2.4. Conditions and Requirements
Given their concerns, participants formulated conditions and requirements for the potential use of AI in QDA, emphasizing tools needed to be transparent, customizable, and supportive of human oversight. To meet privacy considerations, a few participants stressed that AI models should be able to operate in offline mode, ensuring that no data would be uploaded or shared with third parties. Addressing the need for personal involvement and responsibility, several participants emphasized the importance of evaluating the performance and reliability of AI tools before fully integrating them into their workflows. Many participants were particularly concerned about retaining control and agency over the analysis process – with several participants stressing the ability to adjust the level of AI’s influence, preferring a more interactive collaboration with AI rather than a fully automated analysis. A majority of participants also expressed a strong desire for mechanisms to check and verify AI outputs, ensuring the results could be trusted. Several participants further wanted explanations for AI-generated codes: “I would want an explanation for each code, or at least an example from the data to show where the decision was coming from.” Other participants wanted to take random samples of AI outputs and check them or compare them to their own results. Generally, many participants expressed a preference for an interactive approach to AI integration, where they could collaborate with the AI rather than relying on it entirely. Some participants welcomed the possibility to modify AI settings or tune outputs to their needs.
@@
Participants were generally open to AI in QDA, provided it enhanced rather than replaced their skills. They valued AI for offloading repetitive tasks, enabling them to focus on deeper interpretive work. Others saw an opportunity in AI to enrich the analysis with new views and perspectives. Hence, AI was perceived both as a tool for efficiency and as a catalyst for analysis. Participants expressed several concerns and mentioned conditions that had to be met. They suggested developing AI tools that operate offline to allow for greater control over the data’s use and storage. Quality and trustworthiness of AI outputs were another major concern, especially given AI’s current limitations in dealing with nuanced data and the risk of generating false information. Participants stressed the importance of rigorously testing and validating AI tools before integrating them into their research workflows to ensure their reliability in research contexts. Still, they saw potential in AI tools that are developed in alignment with their needs, with a focus on transparency, control, and ethical compliance. Notably, willingness to use AI did not correlate with participants’ experience level, suggesting these conditions are shared across different levels of expertise.
@@ 5. A Framework for AI-based Support that Aligns with HCI Researchers’ Real-Life QDA Workflows
Figure 2. Our framework for integrating AI into QDA, depicting 3 stages from less to more AI involvement, visualizing possible roles for AI we deduced from the data for each stage.:
@@ 5.1. Minimal AI Involvement: Supporting Human-Centric Workflows
AI as an Intelligent User Interface [Uncaptioned image] Figure 3. AI provides step-by-step guidance within QDA software, suggesting functionalities based on user behavior and answering questions to support with problems in tool usage. :
@@
AI as a Productivity Tool [Uncaptioned image] Figure 4. AI generates summary reports of QDA projects, such as identifying the most or least frequent codes, or calculating IRR, allowing the researcher to review these insights and make informed decisions more efficiently. :
@@
AI as a Training Tool [Uncaptioned image] Figure 5. AI supports new researchers in learning and practicing QDA through personalized, interactive instructions, real-time feedback, and simulated collaboration exercises to develop critical thinking and improve coding consistency. :
@@ 5.2. Moderate AI Involvement: Human-AI Collaboration
AI as a Mediator [Uncaptioned image] Figure 6. AI helps researchers resolve coding conflicts and reach consensus efficiently by analyzing discrepancies and suggesting resolutions. :
@@
AI for Validation [Uncaptioned image] Figure 7. AI helps identify errors and inconsistencies while keeping researchers actively engaged, e.g., by suggesting potentially missing annotations and providing them for review. :
@@
AI for Refinement & Explanations [Uncaptioned image] Figure 8. AI provides support by offering concise code definitions and names, e.g., for the final codebook, and suiting quotes for codes. :
@@
AI for Analytics [Uncaptioned image] Figure 9. AI identifies patterns in coding and provides warnings to support decision-making, e.g., when codes overlap or themes could be grouped. :
@@
AI Suggestions [Uncaptioned image] Figure 10. AI analyzes the data and recommends codes which can be reviewed by the researcher and refined by the AI, preserving human agency. :
@@
AI as a Sparring Partner [Uncaptioned image] Figure 11. Researchers and AI collaborate interactively throughout the QDA workflow, from codebook creation, over codebook refinement, to codebook application. :
@@ 5.3. High AI Involvement: AI-Led Workflows
AI with Human-in-the-Loop [Uncaptioned image] Figure 12. AI takes the lead in QDA, e. g., creating a codebook draft (left) or performing initial coding (right), with human oversight throughout. :
@@
AI with Human Approval [Uncaptioned image] Figure 13. AI creates a codebook draft (top) or assigns codes (bottom). A researcher approves before decisions are finalized and changes are applied. :
@@
AI with Conditional Autonomy [Uncaptioned image] Figure 14. AI autonomously handles routine coding but alerts researchers when it encounters data requiring nuanced understanding or deviating from the specified criteria. :
@@ 6. Discussion
6.1. Synthesizing QDA Practices in CSCW
Our findings extend prior discussions of AI-assisted qualitative research by grounding them in a systematic account of how CSCW and HCI researchers actually conduct QDA in practice.
[…]
While AI-supported QDA offers novel opportunities, a major concern regarding the quality of outputs and inherent biases prevails. This is particularly important when analyzing non-English data (Heseltine and Clemm von Hohenberg, 2024; Suter and Meckel, 2024), engaging minority groups, or conducting research in the Global South, as AI is not “culturally agnostic” (Paulus and Marone, 2024) and often privileges dominant or high-frequency representations over others (Feuston and Brubaker, 2021). Perspectives from the Global South and majority world are often marginalized in HCI and CSCW (Kumar and Karusala, 2021; Sturm et al., 2015; Linxen et al., 2021)and in related disciplines (Hasegawa et al., 2024; Sturm et al., 2015). AI-supported QDA risks amplifying this imbalance by embedding dominant interpretative frameworks and overlooking locally relevant meanings (Kumar and Karusala, 2021), particularly when models are trained on Global North or WEIRD datasets (Linxen et al., 2021). These risks are compounded by ongoing challenges in HCI and CSCW around citational justice (Kumar and Karusala, 2021) and transparency of research practices (Wacharamanotham et al., 2020). Addressing them is essential for building culturally sensitive approaches to AI-supported QDA.
@@ 6.2. Putting Our Framework Into Practice
Both prior research and our empirical findings highlight the dual need to respect the interpretive, collaborative nature of QDA while exploring how AI might responsibly augment it. Building on these insights, we propose a nuanced framework for human-AI collaboration in QDA, with varying roles for AI across increasing levels of involvement, tailored to the specific needs of researchers performing QDA. We derive tangible scenarios for AI support that go beyond coding to include tasks such as data preprocessing, onboarding new team members, mediating conflicts, and acting as a conversational partner. We present this framework as flexible rather than prescriptive, recognizing that many factors must be considered when integrating AI into qualitative research. Our findings contribute to ongoing CSCW discussions around responsible, human-centered AI design. They offer timely insights into how HCI researchers are beginning to re-evaluate the role of AI in light of rapidly advancing LLM capabilities.
Our framework aims to empower those who want to incorporate AI tools into QDA, but are unsure of how to approach it or where to intervene. A naive approach is to jump towards AI doing the qualitative coding itself, but this was a contentious topic for our participants. Instead, our framework offers numerous ways AI tools can be incorporated in ways that facilitate QDA without compromising research autonomy. Researchers who are curious about incorporating AI tools such as LLMs but are worried about how this may taint the QDA process can strive for minimal AI involvement, such as AI as a training tool. For companies developing QDA software (e. g., nVivo or MAXQDA), many of which are already adding AI tools to their software, our framework hopes to inspire additional insights into the numerous ways AI tools can be developed to support QDA. Determining the appropriate level and role of AI integration requires careful consideration of factors like research goals, data types, and QDA methods. Below, we outline key considerations for integrating AI into QDA.
@@
Related work shows that coding performance can depend on various factors of complexity, e. g., the required level of judgment and interpretation  (Ortloff et al., 2023; Kirsten et al., 2024). While higher task complexity is an indicator of reduced AI involvement, routine and repetitive tasks like transcription offer the potential for AI leadership, freeing researchers to focus on more complex interpretive tasks. AI should be employed where its capabilities align with the complexity of the task. For example, when the model shows high certainty in a task, it might take a more leading role, while in tasks involving ambiguity or high interpretive judgment, AI could step back. In addition, researchers’ comfort and familiarity with AI tools should guide their level of involvement (Lubars and Tan, 2019; Gauthier and Wallace, 2022). Recent work (Bhatt and Sargeant, 2024; Gauthier and Wallace, 2022; Lam et al., 2024) discusses how restricting AI outputs or displaying performance disclaimers might aid the effective use of AI.
@@
Integrating AI into QDA also involves evaluating ethical considerations, such as data privacy and security, to ensure compliance with ethical and legal standards and regulations. Additionally, participants in human subject studies should be informed during the consent process if AI will be used to analyze their data. Further, the harms of engaging with AI outputs need to be discussed, and safeguards should be applied when feeding data into AI systems. Finally, depending on the level of AI autonomy, a normative discussion on violations of intellectual integrity is necessary. For example, it may be necessary to report the level of AI usage in research outputs to maintain transparency and accountability in the research process. Researchers already use AI throughout their research process (e. g. (Kapania et al., 2024a; Argyle et al., 2023; Hämäläinen et al., 2023; Oppenlaender and Hämäläinen, 2023)), and popular QDA software increasingly integrates AI features – calling the scientific community, and specifically the CSCW and HCI community working at the intersection of humans and technology, to engage this debate and establish guidelines and best practices, e.g., on transparency. Kapania et al. (Kapania et al., 2024a) discuss the ethical guidelines for navigating these risks and concerns.
@@ Appendix B Complete Codebooks from Our Interviews
$\hookrightarrow$ Explainability AI has an explanation for its response “Can they provide justification or explanation?” 7
@@ Appendix C Examples of Participants’ Coding Workflow Drawings
Figure 15. Four examples of coding workflow drawings created by our participants during interviews. The drawings varied considerably in the level of detail they provided. Some included in-depth descriptions of specific actions during each step, while others focused more on visuals, providing details of the process only verbally.: